# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ Abstract
We study strategic interactions in a broker-mediated market. A broker provides liquidity to an informed trader and to noise traders while managing inventory in the lit market. The broker and the informed trader maximise their trading performance while filtering each other’s private information; the trader estimates the broker’s trading activity in the lit market while the broker estimates the informed trader’s private signal. Brokers hold a strategic advantage over traders who rely solely on prices to filter information. We find that information leakage in the client’s trading flow yields an economic value to the broker that is comparable to transaction costs; she speculates profitably and mitigates risk effectively, which, in turn, adversely impacts the informed trader’s performance. In contrast, low signal-to-noise sources, such as prices, result in the broker’s trading performance being indistinguishable from that of a naive strategy that internalises noise flow, externalises informed flow, and offloads inventory at a constant rate.
@@ 1 Introduction
We study a model where a broker (she) provides liquidity to a strategic informed trader (he) and to zero-intelligence noise traders by acting as their counterparty. The broker charges transaction fees to both types of traders. Simultaneously, the broker manages her inventory exposure by trading in the lit market, where she faces both transaction costs and price impact. In this setting, the strategic agents (the broker and the informed trader) operate under partial and heterogeneous information. The informed trader observes a private signal, yet his trading strategy must account for the unobserved broker’s trading activity and its impact on prices, which he estimates from observed market prices. Meanwhile, the broker manages her exposure in the lit market without access to the private signal, which she estimates from her clients’ trading flow or from market price movements.
@@
In this paper, we show how each strategic agent estimates their missing information by filtering relevant observed processes. Specifically, the informed trader relies solely on market prices, which, in a liquid market, have a low signal-to-noise ratio, resulting in low-quality estimates of the broker’s trading speed. In contrast, the broker has access to higher-quality information through her clients’ trading flow, as well as market prices. We show that, in practice, filtering based solely on price information does not produce accurate estimates of the private signal. However, the clients’ trading flow produces significantly better estimates, and can thus improve trading performance and protect the broker’s inventory from adverse selection. We find that the broker holds a strategic advantage over an informed trader who trades exclusively with her; thus, in practice, informed traders need to conceal information or distribute their trading flow across multiple brokers to maintain privacy. [^1] See, for example, Oomen, (2017), for a study of aggregators of multiple brokers.
@@
We employ stochastic control tools in a linear-quadratic framework to derive the optimal adaptive strategies of both agents when they employ prices as the source of information. Both agents maximise their terminal expected wealth subject to inventory penalties. The performance criterion we introduce scales the risk aversion of both agents according to the conditional variance of their respective estimators. We characterise the optimal behaviour of both players and provide conditions for the existence and uniqueness of optimal Markovian strategies, which we derive explicitly. More precisely, given the linear-quadratic structure of the problems, we employ the separation principle and characterise the solutions to the control problems with systems of linear and matrix Riccati differential equations. The matrix Riccati differential equation that characterises the solution to the control problem of the broker is nonstandard. We show that the matrix Riccati differential equation admits a unique solution, which, in turn, ensures a unique optimal Markovian strategy for the broker, when the permanent price impact parameter lies within a certain interval $[0,\delta)$.
We run simulations to study the adaptive strategies of both agents. Our results show that market prices provide poor information for the filtering problems of the informed trader and the broker. Specifically, the broker’s price impact is insufficient to produce a reliable estimate of her trading activity, and the volatility of market prices makes it challenging to infer the informed trader’s private signal accurately. Consequently, the performance of the broker’s strategy in this scenario is akin to that of a naive strategy that externalises the toxic flow, internalises the noise flow, and unwinds inventory based on a trade-weighted average price schedule. In contrast, when the broker employs the trading flow of the informed trader to estimate the signal, her trading performance improves significantly. This improvement arises for two reasons: first, higher-quality signals enable the broker to speculate more effectively on future prices; second, she can manage externalisation and internalisation with better precision. Specifically, the broker increases externalisation and speculative activity when signal values are high or when she is more certain about her estimate.
@@
Our work combines elements of the execution literature for liquidity takers, and the market making literature for liquidity providers. Early works on algorithmic trading include Bertsimas and Lo, (1998), Ho and Stoll, (1983), Almgren and Chriss, (2001), and Avellaneda and Stoikov, (2008), which spurred numerous subsequent studies on execution and market making; see e.g., the textbooks Cartea et al., (2015) and Guéant, (2016). Our work also relates to stochastic games between strategic agents. Bank et al., (2021) develop a continuous-time trading game with two interacting markets: the dealer market where agents trade at an equilibrium price and liquidity providers charge a fee to execute liquidity takers’ demands, and the open market where the demands in the first market are absorbed in an unaffected, martingale price setting. Fu et al., (2022) introduce a portfolio liquidation game with self-exciting order flow that uses Hawkes processes with intensities that depend on past orders. Micheli et al., (2023) study open-loop, closed-loop, and central planning equilibrium of a game in which a interact through the impact they have in prices. Herdegen et al., (2023) investigate market makers competing for market orders of clients in a one-shot Nash competition. Abi Jaber et al., (2023) study a general class of infinite-dimensional linear-quadratic stochastic games with mean-field interaction between the agents. Chen et al., (2023) consider an optimal liquidation model where an investor executes meta-orders which trigger child orders and have both instantaneous and permanent feedback on future order flow. In our paper, the flow of information plays a key role. Here, only the broker has impact in prices because she is the only one that trades in the lit market (where trading is anonymous), and this impact is unknown to the informed trader. On the other hand, the informed trader observes a trading signal that the broker does not observe. Here, both agents (the broker and the informed trader) filter observed processes in search of the information they do not possess.
The combination of filtering techniques and stochastic control tools has been employed in the literature to study algorithmic trading. Campi and Zabaljauregui, (2020) study market making with filtering in a stochastic impulse control setting, Lim and Quenez, (2015) solve a portfolio optimisation problem under partial information for agents with no insider information in a stochastic price jump setting, Allan and Cohen, (2019) formulate a control problem to estimate parameters in the continuous-time Kalman-Bucy filter, Drissi, (2022) proposes models for optimal execution when the drift is estimated from prices, and Allan and Cohen, (2020) use pathwise stochastic control to construct filters that are robust to parameter uncertainty. Recently, Knochenhauer et al., (2024) combine the Kalman-Bucy filter with a linear-quadratic Gaussian control problem to study the problem of costly information acquisition. Pham, (2001) is concerned with mean-variance hedging when the drift of asset prices is partially observable and Pham and Quenez, (2001) search for the optimal portfolio in a stochastic volatility model where investors only observe prices. Aase et al., (2012) investigate the optimal trading of an insider using filtering theory. Firoozi et al., (2018) study a multi-agent (mean-field) game in which the agents attempt to filter a hidden process driving an observable common process. Casgrain and Jaimungal, (2018) characterise optimality in a large stochastic game with heterogeneous agents in the mean-field limit where agents filter latent factors. Finally, Cartea et al., 2024a employ filtering techniques to estimate the fundamental value of an asset from the trading flow of liquidity takers in decentralised trading venues. In this paper, the informed trader and the broker filter distinct information from each other. The broker does not observe the price signal while the informed trader does not observe the permanent price impact from the broker’s activity. As such, we introduce different filtering objects to both of them instead of a common filtering framework.
@@
Closest to our work are Cartea and Sánchez-Betancourt, (2022) and Bergault and Sánchez-Betancourt, (2024). Both papers study versions of a game between a broker and her clients. [^2] There are other strands of the literature that are close to our work, such as Butz and Oomen, (2019) and Cartea et al., (2023) who develop tools to decide when to internalise or externalise in foreign exchange markets. Recently, Barzykin et al., (2024) study filtering of toxic flow in the externalisation of order flow. In the algorithmic trading literature, trading with signals appears in Cartea and Jaimungal, (2016) as the order flow of the market, or more generally in Neuman and Voß, (2022) as a predictable finite-variation process affecting asset prices. In Cartea and Sánchez-Betancourt, (2022), the broker faces an informed and an uninformed trader. The broker and the informed trader optimise their criteria, and they include ambiguity aversion to disentangle their interactions. [^3] Recently, Cartea et al., 2024b solve the perfect information Nash equilibrium of this game. Here we do not assume perfect information; instead, both agents filter the information they do not have. On the other hand, Bergault and Sánchez-Betancourt, (2024) study a large number of informed traders, each of which have a private and a common signal. They characterise, in the mean-field limit, the Nash equilibrium of the trading strategies. In our paper we depart from these approaches. Both the informed trader, and the broker, attempt to filter the missing information in their respective frameworks. More precisely, the broker filters the signal that only the informed trader has, and the informed trader filters the trading speed of the broker, which is only known to the broker and that has permanent impact on prices. As far as we are aware, this is the first time such a problem is studied in the algorithmic trading literature; more precisely, to the best of our knowledge this is the first time that liquidity providers and liquidity takers optimise their continuous-time stochastic control problems while filtering the information that they do not have.
The remainder of the paper proceeds as follows. Section 2 reviews the key elements of the model. Section 3 formulates the filtering problem of the informed trader and solves his stochastic control problem. Section 4 introduces the filtering problem of the broker and solves her stochastic control problem. Here, we prove that there is a range of values that the permanent price impact parameter can take, for which one can show existence and uniqueness of the optimal solution. Section 4.1.2 proposes two alternative filters for how the broker extracts an estimate of the signal that she does not observe. Section 5 carries out simulations to study the performance of the strategies we derive against three benchmarks. For given model parameters, we find evidence that the strategy that externalises the order flow of the informed trader, and internalises the order flow of the uninformed trader while offloading inventory, has a performance that is indistinguishable from the optimal strategy we derive. This is because in our setting, as well as in practice, the signal-to-noise ratio is low, and thus, the filtered signal is not informative. We show that the alternative filters we build have an economic value that is in the order of magnitude as transaction costs. We study the robustness of the best-performing estimate and find that it is robust to inventory misspecification for the earlier parts of the trading day, and it has a poor performance towards the end of the trading day.
2 The Model
In our model, a broker (she) provides liquidity to a strategic informed trader (he) and to uninformed noise traders. This section introduces the general competition setup between the two strategic players; the informed trader and the broker. Section 3 solves the optimal trading problem of the informed trader, and Section 4 solves the optimal trading problem of the broker.
@@ 2 The Model
The broker provides liquidity to the informed trader and to the uninformed trader by acting as counterparty to their trades throughout the trading window ${\mathfrak{T}}$. The informed trader and the uninformed trader carry out their trades at speeds $\eta$ and $\xi$ respectively. Positive values of the trading speed denote buying, and negative values denote selling. As remuneration for this service, the broker charges transaction costs to both traders. In our model, the transaction costs charged by the broker are linear in her clients’ trading speed and are modelled by the positive coefficients $\mathfrak{b}$ for the informed trader, and $\mathfrak{c}$ for the uninformed trader. [^4] If the broker charges the same fee to all clients, then we take $\mathfrak{b}=\mathfrak{c}$. In parallel to providing liquidity, the broker manages her exposure and inventory by trading in the open (or lit) market at a trading speed $\nu$. We assume that the broker incurs linear transaction costs modelled by the positive constant $\mathfrak{a}$, and that her trading activity in the market generates a linear permanent impact modelled by the coefficient $\mathfrak{p}>0$.
@@
where $\kappa^{\alpha}\geq 0$ is the mean-reversion parameter, $\sigma^{\alpha}>0$ is the volatility of the signal, and $W^{\alpha}$ is a Brownian motion. The signal and price processes may share common stochastic factors, so we assume that $W^{\alpha}$ and $W^{S}$ are correlated and write $\langle W^{S},W^{\alpha}\rangle_{t}=\rho\,t$, where $\rho\in[-1,1]$. [^5] See Drissi, (2022) and Bergault et al., (2022) for models of optimal execution and market making with Ornstein–Uhlenbeck dynamics and predictive signals.
@@ 3.1   Learning equations
where $\theta^{B},\sigma^{B}>0$, $\nu_{0}$ is a square-integrable random variable measurable with respect to ${\mathcal{F}}^{I}_{0}$, and $W^{B}$ is a Brownian noise independent of $W^{S}$. The dynamics (3.2) are a key modelling assumption to derive the closed-form solution to the informed trader’s adaptive control problem, i.e., with filtering. The parameters $\theta^{B}$ and $\sigma^{B}$ encode what the informed trader believes about the dynamics of the trading speed of the broker.
[…]
@@
Note that the filtration $({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}}$ is now fixed. This is because the latent $\nu$ is assumed to be exogenous and does not depend on the informed trader’s control. As such, the observation process $Y$ is already decoupled from the informed trader’s control and there is no “chicken and egg” problem as in Section 9.1 of Bensoussan, (2018).
3.2   The model
@@
[…]
@@
where $\beta_{0},\beta_{1},\rho_{0},\rho_{1}>0$, $\mathfrak{x}=(s,x^{I},q^{I},\hat{\nu})$, and the expectation is under $\mathbb{P}^{I}$. The first two terms in the expectation in (3.7) are the informed trader’s terminal wealth composed of the marked-to-market value of his inventory and his terminal cash. The last two terms represent a terminal and running risk aversion components, respectively. In particular, we introduce the new terms $\beta_{1}\,\mathbb{V}^{I}$ and $\rho_{1}\,\mathbb{V}^{I}$ to adjust risk aversion according to the conditional variance $\mathbb{V}^{I}_{t}$. [^7] When $\beta_{1}=\rho_{1}=0$ we recover a performance criterion that is widely used in algorithmic trading; see e.g., Donnelly, (2022) for a recent review article. These terms penalise any terminal or running inventory according to the informed trader’s uncertainty because higher conditional variance implies more uncertainty in the estimation of the true $\nu$. The associated control problem is
@@
3.3   Hamilton–Jacobi–Bellman equation
00 0 $=\partial_{t}H^{I}-(\rho_{0}+\rho_{1}\,\mathbb{V}^{I}_{t})\left(q^{I}\right)^{2}+\sup_{\eta}\big{[}(\mathfrak{p}\,\hat{\nu}+\alpha)\,\partial_{s}H^{I}-\eta\left(s+\mathfrak{b}\,\eta\right)\,\partial_{x^{I}}H^{I}+\eta\,\partial_{q^{I}}H^{I}\big{]}$
$\qquad+\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\partial_{s\hat{\nu}}H^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\alpha}H^{I}+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\hat{\nu}}H^{I}$ (3.9)
$\qquad+\sigma^{S}\,\sigma^{\alpha}\,\rho\,\partial_{s\alpha}H^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\mathbb{V}^{I}_{t}\,\rho}{\sigma^{S}}\,\partial_{\alpha\hat{\nu}}H^{I}\,,$ (3.10)
with terminal condition
The supremum in the Hamiltonian in (3.10) is attained at
$\eta^{*}=\frac{\partial_{q^{I}}H^{I}-\partial_{x^{I}}H^{I}\,s}{2\,\partial_{x^{I}}H^{I}\,\mathfrak{b}}\,.$ (3.11)
After substitution, employ the ansatz
$H^{I}=x^{I}+h^{I}(t,s,q^{I},\alpha,\hat{\nu})\,$
to obtain that $h^{I}$ satisfies the partial differential equation (PDE)
with terminal condition
Next, use the ansatz
to obtain the system of differential equations
00 0 $=-(\rho_{0}+\rho_{1}\mathbb{V}^{I}_{t})+(g_{2}^{I})^{\prime}+\frac{(g_{2}^{I})^{2}}{\mathfrak{b}}$ (3.12)
00 0 $=\partial_{t}\,g_{1}^{I}+(\mathfrak{p}\hat{\nu}+\alpha)+\frac{g_{1}^{I}\,g_{2}^{I}}{2\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{1}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{1}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{1}^{I}\,,$
[…]
00 0 $=\partial_{t}\,g_{0}^{I}+\frac{(g_{1}^{I})^{2}}{4\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{0}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{0}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{0}^{I}$
[…]
with terminal conditions
$g_{2}^{I}(T)\quad$ $=-(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{T})\,,$ (3.15)
$g_{1}^{I}(T,\alpha,\hat{\nu})$ $=\qquad 0\,,\qquad$ (3.16)
$g_{0}^{I}(T,\alpha,\hat{\nu})$ $=\qquad 0\,.\qquad$ (3.17)
The following lemma solves the system of PDEs (3.12)-(3.13)-(3.14).
Lemma 1.
$g_{1}^{I}(t,\alpha,\hat{\nu})$ $:=z_{1}^{I}(t)\,\alpha+z_{2}^{I}(t)\,\hat{\nu}\,,$ (3.18)
$g_{0}^{I}(t,\alpha,\hat{\nu})$ $:=z_{3}^{I}(t)+z_{4}(t)\,\alpha+z_{5}^{I}(t)\,\hat{\nu}+z_{6}^{I}(t)\,\alpha\,\hat{\nu}+z_{7}^{I}(t)\,\alpha^{2}+z_{8}^{I}(t)\,\hat{\nu}^{2}\,,$ (3.19)
solve the system of differential equations (3.14) with terminal conditions (3.15)-(3.16)-(3.17).
Proof.
Observe that (3.12) is a one-dimensional Riccati equation with quadratic term $0{1}/{\mathfrak{b}}>$ 0, “constant” term $-(\rho_{0}^{I}+\rho_{1}^{I}\,\mathbb{V}^{I}_{t})<0$, and terminal value $-(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{t})<0$. Use Lemma 4.1 in Reid, (1972) to obtain that the equation has a unique solution in $(-\infty,T]$, and that it is negative.
Next, to solve (3.13) and (3.14), one may work backwards by substituting the desired form of $g_{0}$ and $g_{1}$ into the system of equations to obtain
00 0 $=\alpha$ $\left[1-\kappa^{\alpha}\,z_{1}^{I}+\frac{g_{2}^{I}\,z_{1}^{I}}{2\,\mathfrak{b}}+(z_{1}^{I})^{\prime}\right]$ (3.20)
[…]
00 0 $=$ $\left[\frac{\mathfrak{p}\,\sigma^{\alpha}\,\rho\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\,z_{6}^{I}+(\sigma^{\alpha})^{2}\,z_{7}^{I}+\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}z_{8}^{I}+(z_{3}^{I})^{\prime}\right]$
[…]
$(z_{1}^{I},z_{2}^{I})\to(z_{6}^{I},z_{7}^{I},z_{8}^{I})\to(z_{4}^{I},z_{5}^{I})\to z_{3}^{I}\,.$ (3.23)
In the above order, the ODEs are linear, so there exist $\{z_{i}^{I}\}_{i=1,\cdots,8}$ satisfying (3.13) and (3.14). ∎

### v2

@@ Abstract
We study strategic interactions in a broker-mediated market in which agents learn and exploit each other’s private information. A broker provides liquidity to an informed trader and to noise traders while managing inventory in a lit market. The informed trader infers the broker’s trading activity in the lit market, while the broker estimates the trader’s private signal. Information leakage in the client’s trading flow generates economic value for the broker that is comparable in magnitude to transaction costs: the broker can speculate profitably and manage risk more effectively, which in turn adversely affects the informed trader’s performance. Brokers therefore hold a strategic advantage over traders who rely solely on prices to filter information. When the broker only relies on prices rather than client trading flow to infer information, their trading performance becomes indistinguishable from the performance of a naive strategy that internalises noise flow, externalises informed flow, and offloads inventory at a constant rate.
Keywords: algorithmic trading, externalisation, internalisation, filtering, signal-to-noise, brokers.
@@ 1 Introduction
We study a model in which a broker (“she”) provides liquidity to an informed trader (“he”) and to noise traders by acting as their counterparty. Simultaneously, the broker manages her inventory exposure by trading in the lit market, where she faces both transaction costs and price impact. In this setting, the strategic agents, i.e., the broker and the informed trader, operate under partial and heterogeneous information. The informed trader observes a private signal, and his trading strategy accounts for the price impact of the broker’s unobserved trading activity, which he estimates from market prices. Meanwhile, the broker manages her exposure in the lit market, and her strategy accounts for the unobserved trader’s private signal, which she estimates either from her clients’ trading flow or from market prices.
@@
In this paper, we show how each agent estimates the information they do not observe by filtering relevant processes. Specifically, the informed trader relies solely on market prices, which in liquid markets have a low signal-to-noise ratio, resulting in imprecise estimates. We show that, in practice, filtering based only on prices does not yield accurate estimates of the broker’s activity. In contrast, the broker has access to higher-quality information through her clients’ trading flow, in addition to market prices. The trading flow of informed clients produces significantly more informative signals, which improves the broker’s trading performance and protects her against adverse selection. As a result, the broker holds a strategic advantage over an informed trader who trades exclusively with her. In practice, informed traders therefore need to conceal information or distribute their trading flow across multiple brokers to preserve privacy. [^1] See Oomen (2017) for a study of broker aggregation.
@@
We employ stochastic control tools to derive the optimal adaptive strategies of both agents when prices are used as the information source. Both agents maximise terminal expected wealth subject to inventory penalties, with risk aversion scaled according to the conditional variance of their respective estimators. We characterise the optimal behaviour of both players and provide conditions for the existence and uniqueness of optimal Markovian strategies, which we derive explicitly. Owing to the linear–quadratic structure of the problem, we apply the separation principle and characterise the solutions as systems of linear and matrix Riccati differential equations (RDEs). The matrix RDE governing the broker’s problem is nonstandard. When the broker learns the signal from prices, we show that this RDE admits a unique solution, which in turn guarantees a unique optimal Markovian strategy. When the broker instead uses the informed trader’s order flow, the matrix RDE admits a unique solution up to a non-explosion condition.
We illustrate the adaptive strategies of both agents in extensive numerical experiments. Our results confirm that market prices constitute a poor information source for filtering: the broker’s price impact is too small relative to price noise to produce reliable estimates of her own trading activity, and price volatility makes it difficult to infer the informed trader’s private signal accurately. Consequently, the broker’s performance in this case resembles that of a naïve strategy that internalises noise flow, externalises toxic flow, and unwinds inventory at a constant rate.
In contrast, when the broker uses the informed trader’s order flow to estimate the signal, her performance improves substantially. This improvement arises for two reasons: higher-quality signals allow the broker to speculate more effectively on future prices, and they enable more precise internalisation and externalisation decisions. In particular, the broker increases externalisation and speculative trading when signal values are high or when her estimates are more precise. We show that the economic value of the improved filtering is of the same order of magnitude as transaction costs. Finally, we study the robustness of the best-performing estimator and find that it remains robust to inventory misspecification during the early part of the trading horizon.
@@
Our work combines elements from the execution literature for liquidity takers and the market-making literature for liquidity providers. [^2] Early contributions to algorithmic trading include Bertsimas and Lo (1998), Ho and Stoll (1981), Almgren and Chriss (2001), and Avellaneda and Stoikov (2008), which spurred a large body of subsequent research on execution and market making; see, for example, the textbooks Cartea et al. (2015) and Guéant (2016). Our work also relates to the literature on stochastic games. A closely related contribution is Bank et al. (2021), which develops a continuous-time trading game with two interacting markets: a dealer market in which agents trade at an equilibrium price and liquidity providers charge a fee to execute liquidity takers’ demands, and an open market in which the demands originating in the dealer market are absorbed at an unaffected martingale price. [^3] A broader literature studies stochastic games in financial markets. Examples include portfolio liquidation games with self-exciting order flow Fu et al. (2022), games with price-impact interaction under open-loop, closed-loop, and centralised equilibria Micheli et al. (2023), and Nash competition among market makers for client order flow Herdegen et al. (2023). Related work analyses infinite-dimensional games with mean-field interactions Abi Jaber et al. (2023), as well as optimal liquidation problems with feedback effects on future order flow Chen et al. (2023). In contrast, the flow of information plays a central role in our setting. Only the broker has price impact, as she is the sole agent trading in the lit market where transactions are anonymous, and this impact is unobserved by the informed trader. Conversely, the informed trader observes a private trading signal that is not available to the broker. As a result, the broker and the informed trader engage in filtering observed processes to infer the information they do not possess. [^4] Our work also relates to literature combining filtering techniques and stochastic optimal control. Early contributions include hedging (Pham (2001); Pham and Quenez (2001)), insider trading models (Aase et al. (2012)), and portfolio choice (Lim and Quenez (2015)). Filtering is also studied in mean-field games with heterogeneous agents (Casgrain and Jaimungal (2018); Firoozi et al. (2018)), robust control (Allan and Cohen (2019, 2020)), market making (Campi and Zabaljauregui (2020); Drissi (2022); Cartea et al. (2024)), and costly information acquisition (Knochenhauer et al. (2024)).
@@
Closest to our work are Cartea and Sánchez-Betancourt (2025) and Bergault and Sánchez-Betancourt (2025), both of which study strategic interactions between a broker and her clients. In Cartea and Sánchez-Betancourt (2025), the broker faces both informed and uninformed traders, and both the broker and the informed trader optimise their objectives under ambiguity aversion to disentangle their strategic interactions. [^5] Related strands of the literature include work on internalisation and externalisation decisions in FX markets Butz and Oomen (2019); Barzykin et al. (2023); Cartea et al. (2023), filtering of toxic flow Barzykin et al. (2024), and trading with predictive signals in execution problems Cartea and Jaimungal (2016); Neuman and Voß (2022); Drissi (2022); Bergault et al. (2022). Recently, Cartea et al. (2025) solve the perfect-information Nash equilibrium of the game. In contrast, we do not assume perfect information; instead, both agents must filter and infer the information they do not observe. On the other hand, Bergault and Sánchez-Betancourt (2025) study a large population of traders with both private and common signals, and characterise the Nash equilibrium in the mean-field limit. We depart from this approach by modelling asymmetric information acquisition. In our framework, both agents actively filter the information they do not observe within their respective decision problems. To the best of our knowledge, this paper, together with Wu and Jaimungal (2024), is among the first to study settings in which both liquidity providers and takers optimise their strategies while filtering missing information. While Wu and Jaimungal (2024) analyse the Nash equilibrium in a multidimensional setting, our approach adopts a Stackelberg-type formulation.
The remainder of the paper proceeds as follows. Section 2 reviews the key elements of the model. Section 3 formulates the filtering problem of the informed trader and solves his stochastic control problem. Section 4 introduces the filtering problems of the broker and solves both of her stochastic control problems. In particular, we propose two alternative filters (and solve their respective control problems) for how the broker extracts an estimate of the signal that she does not observe. Section 5 carries out simulations to study the performance of the strategies we derive against three benchmarks.
2 Modelling framework
@@
The broker provides liquidity to the informed trader and to the uninformed trader by acting as counterparty to their trades throughout the trading window $\mathfrak{T}$}. The informed trader and the uninformed trader carry out their trades at speeds $\eta$ and $\xi$ respectively. Positive values of the trading speed denote buying, and negative values denote selling. As remuneration for this service, the broker charges transaction costs to both traders. In our model, the transaction costs charged by the broker are linear in her clients’ trading speed and are modelled by the positive coefficients $\mathfrak{b}$ (for the informed trader), and $\mathfrak{c}$ (for the uninformed trader). [^6] If the broker charges the same fee to all clients, then we take $\mathfrak{b}=\mathfrak{c}$. In parallel to providing liquidity, the broker manages her exposure and inventory by trading in the open (or lit) market at a trading speed $\nu$. We assume that the broker incurs linear transaction costs modelled by the positive constant $\mathfrak{a}$, and that her trading activity in the market generates a linear permanent impact modelled by the coefficient $\mathfrak{p}>0$.
@@
where $\kappa^{\alpha}\geq 0$ is the mean-reversion parameter, $\sigma^{\alpha}>0$ is the volatility of the signal, and $W^{\alpha}$ is a Brownian motion. The signal and price processes may share common stochastic factors, so we assume that $W^{\alpha}$ and $W^{S}$ are correlated and write $\langle W^{S},W^{\alpha}\rangle_{t}=\rho\,t$, where $\rho\in[-1,1]$.
@@ 3.1  Learning equations
where $\theta^{B},\sigma^{B}>0$, and $W^{B}$ is a Brownian noise independent of $W^{S}$. The dynamics in (3.2) are a key modelling assumption to derive the closed-form solution to the informed trader’s adaptive control problem, i.e., with filtering. The parameters $\theta^{B}$ and $\sigma^{B}$ encode what the informed trader believes about the dynamics of the trading speed of the broker.
Here, we use the Kalman-Bucy filter for processes with linear drift and diffusion terms (see e.g., Bain and Crisan (2009) and Chapter 22 in Cohen and Elliott (2015)) to solve the informed trader’s stochastic filtering problem of $\nu$ from observing $Y$. The filter is characterised by the dynamics of its mean and (co)variance. More precisely, assume (3.2) holds and let ${\mathcal{F}}^{I}_{t}=\sigma(\alpha_{s},Y_{s}:\,s\in[0,t])\vee{\mathcal{N}}$ where $\mathcal{N}$} is the set of $\mathbb{P}$-null sets of the full sigma-algebra $\mathcal{F}$}, ${\mathcal{F}}^{I}=\sigma\left(\bigcup_{t\in{\mathfrak{T}}}{\mathcal{F}}^{I}_{t}\right)$, and $\mathcal{C}\vee\mathcal{D}$ denotes the smallest sigma-algebra containing $\mathcal{C}$ and $\mathcal{D}$. Let $\hat{\nu}_{t}=\mathbb{E}\left[\nu_{t}\,|\,{\mathcal{F}}^{I}_{t}\right]$ be the projection of $\nu_{t}$ onto ${\mathcal{F}}^{I}_{t}$ and let $\mathbb{V}^{I}_{t}=\mathbb{E}\left[(\nu_{t}-\hat{\nu}_{t})^{2}\,|\,{\mathcal{F}}^{I}_{t}\right]$ be its conditional variance. The Kalman and Bucy (1961) filter gives
@@
Note that the filtration $({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}}$ is now fixed. This is because the latent $\nu$ is assumed to be exogenous and does not depend on the informed trader’s control. As such, the observation process $Y$ is already decoupled from the informed trader’s control and there is no “chicken and egg” problem as in Section 9.1 of Bensoussan (2018).
3.2  The control problem of the informed trader
@@
${\mathcal{A}}^{I}:=\left\{\eta=\left(\eta_{t}\right)_{t\in{\mathfrak{T}}}\,:\,\eta\text{ is $({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}}-$progressively measurable }\text{and}\,\,\mathbb{E}^{I}\left[\int_{0}^{T}(\eta_{s})^{2}\,ds\right]<\infty\right\}\,$, (3.6)
@@
[…]